At 25 °C, the vapor pressure of acetone is 230 mmHg and the vapor pressure of toluene is 284.5 mmHg. What is the total vapor pressure at 25 °C above an equimolar mixture of these two miscible liquids?
1
Expert's answer
2015-11-02T12:12:30-0500
Total vapor pressure = partial pressure (acetone) + partial pressure (toluene)
Partial pressure = molfraction * vapor pressure
Molfraction (acetone) = molfraction (toluene) = 0.5 (only for equimolar substances)
Partial pressure (acetone) = 0.5 * 230mmHg = 115mmHg

Partial pressure (toluene) = 0.5 * 284.5mmHg = 142.25mmHg

Total vapor pressure = 115mmHg + 142.25mmHg = 257.25mmHg